OpenStudy (gamereternity):

If f(x) varies directly with x, and f(x) = 72 when x = 8, find the value of f(x) when x = 3. A. 9 B. 18 C.<=my answer 27 D. 192

OpenStudy (gamereternity):

@MaddieLB

OpenStudy (photon336):

you're correct :) \[y = k*x\] \[\frac{ y }{ x } = k => \frac{ 72 }{ 8 } = 9 = k \] \[f(x) = 9x\]

OpenStudy (gamereternity):

@Photon336 but i chose c which is 27 so is it that or a 9

OpenStudy (photon336):

well @gamereternity 9*x is the function. when we plug in x = 3 we get 27 so you're correct.

OpenStudy (gamereternity):

ooh ok ty very much